OpenStudy (anonymous):

the diameter of Z is 5in. what is its area in terms of pie

OpenStudy (abdullah1995):

pi*2.5^2

OpenStudy (akashdeepdeb):

Area of a circle (if that's what you have to find) = \(\pi * r^2\) We also know that, \(d = 2r ~~\therefore r = \frac{d}{2}\) Now, because we know 'd', we substitute this back into the area equation, like so:\[Area ~~= ~~\pi * (\frac{d}{2})^2\] d = 5 inches. Can you find the area now? :)